Opened 10 years ago

Last modified 3 years ago

#274 assigned flight processing

EU09/02, flight day 239a/2009, Swiss alps

Reported by: mark1 Owned by: knpa
Priority: alpha 4 medium Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2009/flight_data/other/EU09_02-2009_239a_Swiss_alps

Data arrived from ARSF via SATA disk L on 7th October 2009.

Scientific objective: Mapping glacier debris cover thickness and extent using aerial photography, LiDAR and spectral emissivity

Priority: alpha-4-medium

PI: David Rippin

Sensors:

  • Eagle (18/05/10)
  • Hawk (18/05/10)

Change History (16)

comment:1 Changed 10 years ago by mark1

Still awaiting logsheet and rinex

comment:2 Changed 10 years ago by mark1

  • Description modified (diff)

No LIDAR for this flight - initially an ALS60 was borrowed but due to conditions and problems with the sensor collection was abandoned.

comment:3 Changed 10 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Started hyperspectral processing.

comment:4 Changed 10 years ago by mark1

applanix processing has been done, using the GPS data collected for GB07-09/10 239b.

comment:5 Changed 10 years ago by knpa

Generated config file.

comment:6 Changed 10 years ago by knpa

no sync on hawk 1 and 2.

comment:7 Changed 10 years ago by knpa

SCTs

e h
1 0.00 1.36
2 0.00 1.27

comment:8 Changed 10 years ago by knpa

HS delivery prepared, awaiting delivery check.

comment:9 Changed 10 years ago by mark1

Hyperspectral delivery checked. Made edits to read me file, changed permissions on lev1 files to make readable. Viewed data in fastQC and looks good.

Cannot get example command to work on any of the lev1 files - maybe HDF files are corrupt? Command works fine on other data but not this set, might need reprocessing.

comment:10 Changed 10 years ago by mark1

Checked the re-processed level1 files and they look good and example command works.

comment:11 Changed 10 years ago by knpa

Hyperspectral delivered today (18/05/10)

comment:12 Changed 10 years ago by knpa

Repository is full, need to rsync when there is room.

comment:13 Changed 10 years ago by knpa

  • Description modified (diff)

comment:14 Changed 9 years ago by knpa

  • Component changed from Processing: general to Archiving

project r-synced.

comment:15 Changed 6 years ago by tipo

Reprocessed hyperspectral.

comment:16 Changed 3 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.